# Oluwanifemi Bamgbose§, Akshay Kalkunte§, Sai Rajeswar§, Sathwik Tejaswi Madhusudhan§ § ServiceNow

### v1

@@ Abstract
Large Audio Language Models (LALMs) are rapidly advancing, but evaluating them remains challenging due to inefficient toolkits that limit fair comparison and systematic assessment. Current frameworks suffer from three critical issues: slow processing that bottlenecks large-scale studies, inconsistent prompting that hurts reproducibility, and narrow task coverage that misses important audio reasoning capabilities. We introduce LALM-Eval, an efficient and comprehensive evaluation framework for LALMs. Our system achieves a speedup of up to 127% over existing toolkits through optimized batch processing and parallel execution, enabling large-scale evaluations previously impractical. We provide standardized prompting protocols and flexible configurations for fair model comparison across diverse scenarios. Additionally, we introduce two new evaluation categories: LLM-Adaptive Diarization for temporal audio understanding and Spoken Language Reasoning for complex audio-based cognitive tasks. Through evaluation across 380+ tasks, we reveal significant gaps in current LALMs, particularly in temporal understanding and complex spoken language reasoning tasks. Our findings also highlight a lack of standardization in instruction modality existent across audio benchmarks, which can lead up performance differences up to 9.5 absolute points on the challenging complex instruction following downstream tasks. LALM-Eval provides both practical evaluation tools and insights into model limitations, advancing systematic LALM development.
GitHub Repository: https://github.com/ServiceNow/LALM-Eval
@@ 3 LALM Evaluation Challenges
LALM-Eval 3.6 ($\downarrow$48.75%) 3.65 ($\uparrow$ 95.19%)
@@ 3.3 Comprehensive Task Category Coverage
Figure 1: Architecture overview of LALM-Eval evaluation framework. Our system comprises three core components: (1) Config module for hierarchical task configuration and standardized prompting, (2) Request Controller managing token-based concurrency limits across all engines with adaptive retry mechanisms, and (3) Concurrent Engines executing parallel model evaluation with dataset sharding. The Request Controller maintains a global token pool accessible to all engines, enabling efficient resource utilization and scalable throughput. Multiple concurrent connections between the controller and inference models illustrate parallel request dispatch, with each engine supporting multi-model evaluation on targeted datasets.
Figure 2: Task distribution and coverage in LALM-Eval. Our framework encompasses six major task categories with balanced representation: Speech Recognition (ASR variants), Paralinguistics (emotion, speaker, accent recognition), Spoken Language Understanding (QA, translation, summarization), Audio Understanding (scene, music), Spoken Language Reasoning (function calling, coding, instruction following), and Safety & Security (robustness, spoofing detection). The distribution reflects comprehensive coverage from basic perception to complex reasoning, with novel emphasis on prompted temporal understanding and audio-conditioned cognitive tasks.
4 LALM-Eval
In response to the challenges existing in current audio understanding evaluation toolkits, we propose our standardized, efficient, highly customizable evaluation framework, LALM-Eval, as detailed in Figure 1.
Our LALM-Eval is composed of 3 primary components, including: Config, Request Controller and Concurrent Engines. The Config module defines a structured and hierarchical representation of customizable configurations, enabling flexible and transparent evaluation settings. The Request Controller is responsible for managing token requests and coordinating execution across the framework. Finally, the Concurrent Engines module carries out task-specific evaluations in parallel, where each engine can support multi-model evaluation tailored to particular tasks. In the following sections, we introduce our architecture design in more detail to address the aforementioned challenges in Section 3.
@@ 4.1 Inference Efficiency
As illustrated in Figure 1, our proposed LALM-Eval maximizes inference efficiency through a token-based request scheduling architecture. More specifically, we introduce a Central Request Controller that maintains and regulates a pool of available tokens accessible to all models across all evaluation engines. Here, a token refers to a concurrency slot representing permission to issue one inference request (not a model input token), which is acquired before dispatch and released upon completion. Each concurrent engine-specific requester periodically draws from the global pool. Within each of these engines, multiple models are executed concurrently on a targeted dataset, with inference calls dispatched in parallel to fully exploit available computational resources. This architecture ensures that evaluation throughput is not bottlenecked by model or engine-specific constraints, but rather governed solely by user-defined request limits set globally, providing both scalability and predictable performance guarantees. Furthermore, we allow user-specified retry counts on request errors, enabling users to set higher request limits with the assurance that occasional failures will be re-tried and successfully completed, thereby offering a tunable balance between throughput and reliability.
Furthermore, our LALM-Eval implements a layered request synchronization strategy that adaptively staggers request wait times across concurrent models. This design increases the probability that all models processing a given dataset segment complete their inference in a temporally aligned manner. By reducing discrepancies in model response times, the strategy minimizes idle periods within each engine, thereby mitigating intra-engine waiting time and improving overall throughput efficiency.
@@ 4.2 Customizable Evaluation Configurations
LALM-Eval is highly customizable with structured task coverage as presented in Figure 2, from dataset usage to inference to evaluations.
@@
Great effort has been made into making LALM-Eval as customizable as possible, while still being intuitive to use. First, any number of open and closed source models, across any number of datasets, can be run. Each model can have its own specific temperature and max-token settings, which override (also customizable) task-specific temperature settings. Each model endpoint also has specific run specifications that can be changed, such as concurrent allowed requests, error retry limit, timeout before retry, and audio chunk size. This is all in order to maxmize resource utilization and minimize evaluation time.
@@
LALM-Eval framework is also designed for granular control over evaluation steps by allowing for customizable metric assignment on a per-dataset and per-task basis. For instance, LLM-as-judge supports configurable concurrency to maximize the throughput for evaluation stage. For a more comprehensive understanding of model performance, the framework offers configurable aggregation metrics. This capability allows for the multi-dimensional analysis of task and metric results, providing a holistic view that extends beyond simple, individual scores or sub-tasks.
@@ 5 Results & Discussion
Empirical evaluations across all task categories using our proposed LALM-Eval are provided in Table 4. We use GPT-4o-mini as judge for LLM-judge metrics due to its advanced capability.
@@ 5.1 Inference Efficiency
We perform empirical evaluation to compare LALM-Eval against existing evaluation kits: AudioBench Wang et al. (2025a), VoiceBench Chen et al. (2024), and Kimi-Eval Ding et al. (2025). Our analysis focuses on the two key metrics RTF and Processed Samples per Second detailed in Section 3.1. We leverage 500 audio samples from 3 diverse datasets: librispeech-clean-test, ClothoAQA, and MELD-Emotion. The evaluation is conducted on three different LALMs, including: Qwen2.5-Omni-7B Xu et al. (2025), Phi-4-Multimodal Abouelenin et al. (2025) and Voxtral-Mini-3B Liu et al. (2025). For conciseness, we report the averaged metric across all 3 LALMs. More specific details are further provided in Table 5.
@@
As shown in Figure 4, LALM-Eval consistently outperforms existing evaluation kits across all runtime scenarios in two key efficiency metrics. More specifically, our LALM-Eval achieves up to a 127% improvement in Processed Sample per Second and 59% reduction in RTFs compared to the next most competitive evaluation frameworks. More importantly, our Parallel runtime, illustrated in Figure 5, is significantly more efficient than competing frameworks. These empirical results validate our framework as a highly efficient tool for LALM evaluation.

### v2

@@ Abstract
Large Audio Language Models (LALMs) are rapidly advancing, but evaluating them remains challenging due to inefficient toolkits that limit fair comparison and systematic assessment. Current frameworks suffer from three critical issues: slow processing that bottlenecks large-scale studies, inconsistent prompting that hurts reproducibility, and narrow task coverage that misses important audio reasoning capabilities. We introduce AU-Harness, an efficient and comprehensive evaluation framework for LALMs. Our system achieves a speedup of up to 127% over existing toolkits through optimized batch processing and parallel execution, enabling large-scale evaluations previously impractical. We provide standardized prompting protocols and flexible configurations for fair model comparison across diverse scenarios. Additionally, we introduce two new evaluation categories: LLM-Adaptive Diarization for temporal audio understanding and Spoken Language Reasoning for complex audio-based cognitive tasks. Through evaluation across 380+ tasks, we reveal significant gaps in current LALMs, particularly in temporal understanding and complex spoken language reasoning tasks. Our findings also highlight a lack of standardization in instruction modality existent across audio benchmarks, which can lead up performance differences up to 9.5 absolute points on the challenging complex instruction following downstream tasks. AU-Harness provides both practical evaluation tools and insights into model limitations, advancing systematic LALM development.
GitHub Repository: https://github.com/ServiceNow/AU-Harness
@@ 3 LALM Evaluation Challenges
AU-Harness 3.6 ($\downarrow$48.75%) 3.65 ($\uparrow$ 95.19%)
@@ 3.3 Comprehensive Task Category Coverage
Figure 1: Architecture overview of AU-Harness evaluation framework. Our system comprises three core components: (1) Config module for hierarchical task configuration and standardized prompting, (2) Request Controller managing token-based concurrency limits across all engines with adaptive retry mechanisms, and (3) Concurrent Engines executing parallel model evaluation with dataset sharding. The Request Controller maintains a global token pool accessible to all engines, enabling efficient resource utilization and scalable throughput. Multiple concurrent connections between the controller and inference models illustrate parallel request dispatch, with each engine supporting multi-model evaluation on targeted datasets.
Figure 2: Task distribution and coverage in AU-Harness. Our framework encompasses six major task categories with balanced representation: Speech Recognition (ASR variants), Paralinguistics (emotion, speaker, accent recognition), Spoken Language Understanding (QA, translation, summarization), Audio Understanding (scene, music), Spoken Language Reasoning (function calling, coding, instruction following), and Safety & Security (robustness, spoofing detection). The distribution reflects comprehensive coverage from basic perception to complex reasoning, with novel emphasis on prompted temporal understanding and audio-conditioned cognitive tasks.
4 AU-Harness
In response to the challenges existing in current audio understanding evaluation toolkits, we propose our standardized, efficient, highly customizable evaluation framework, AU-Harness, as detailed in Figure 1.
Our AU-Harness is composed of 3 primary components, including: Config, Request Controller and Concurrent Engines. The Config module defines a structured and hierarchical representation of customizable configurations, enabling flexible and transparent evaluation settings. The Request Controller is responsible for managing token requests and coordinating execution across the framework. Finally, the Concurrent Engines module carries out task-specific evaluations in parallel, where each engine can support multi-model evaluation tailored to particular tasks. In the following sections, we introduce our architecture design in more detail to address the aforementioned challenges in Section 3.
@@ 4.1 Inference Efficiency
As illustrated in Figure 1, our proposed AU-Harness maximizes inference efficiency through a token-based request scheduling architecture. More specifically, we introduce a Central Request Controller that maintains and regulates a pool of available tokens accessible to all models across all evaluation engines. Here, a token refers to a concurrency slot representing permission to issue one inference request (not a model input token), which is acquired before dispatch and released upon completion. Each concurrent engine-specific requester periodically draws from the global pool. Within each of these engines, multiple models are executed concurrently on a targeted dataset, with inference calls dispatched in parallel to fully exploit available computational resources. This architecture ensures that evaluation throughput is not bottlenecked by model or engine-specific constraints, but rather governed solely by user-defined request limits set globally, providing both scalability and predictable performance guarantees. Furthermore, we allow user-specified retry counts on request errors, enabling users to set higher request limits with the assurance that occasional failures will be re-tried and successfully completed, thereby offering a tunable balance between throughput and reliability.
Furthermore, our AU-Harness implements a layered request synchronization strategy that adaptively staggers request wait times across concurrent models. This design increases the probability that all models processing a given dataset segment complete their inference in a temporally aligned manner. By reducing discrepancies in model response times, the strategy minimizes idle periods within each engine, thereby mitigating intra-engine waiting time and improving overall throughput efficiency.
@@ 4.2 Customizable Evaluation Configurations
AU-Harness is highly customizable with structured task coverage as presented in Figure 2, from dataset usage to inference to evaluations.
@@
Great effort has been made into making AU-Harness as customizable as possible, while still being intuitive to use. First, any number of open and closed source models, across any number of datasets, can be run. Each model can have its own specific temperature and max-token settings, which override (also customizable) task-specific temperature settings. Each model endpoint also has specific run specifications that can be changed, such as concurrent allowed requests, error retry limit, timeout before retry, and audio chunk size. This is all in order to maxmize resource utilization and minimize evaluation time.
@@
AU-Harness framework is also designed for granular control over evaluation steps by allowing for customizable metric assignment on a per-dataset and per-task basis. For instance, LLM-as-judge supports configurable concurrency to maximize the throughput for evaluation stage. For a more comprehensive understanding of model performance, the framework offers configurable aggregation metrics. This capability allows for the multi-dimensional analysis of task and metric results, providing a holistic view that extends beyond simple, individual scores or sub-tasks.
@@ 5 Results & Discussion
Empirical evaluations across all task categories using our proposed AU-Harness are provided in Table 4. We use GPT-4o-mini as judge for LLM-judge metrics due to its advanced capability.
@@ 5.1 Inference Efficiency
We perform empirical evaluation to compare AU-Harness against existing evaluation kits: AudioBench Wang et al. (2025a), VoiceBench Chen et al. (2024), and Kimi-Eval Ding et al. (2025). Our analysis focuses on the two key metrics RTF and Processed Samples per Second detailed in Section 3.1. We leverage 500 audio samples from 3 diverse datasets: librispeech-clean-test, ClothoAQA, and MELD-Emotion. The evaluation is conducted on three different LALMs, including: Qwen2.5-Omni-7B Xu et al. (2025), Phi-4-Multimodal Abouelenin et al. (2025) and Voxtral-Mini-3B Liu et al. (2025). For conciseness, we report the averaged metric across all 3 LALMs. More specific details are further provided in Table 5.
@@
As shown in Figure 4, AU-Harness consistently outperforms existing evaluation kits across all runtime scenarios in two key efficiency metrics. More specifically, our AU-Harness achieves up to a 127% improvement in Processed Samples per Second and 59% reduction in RTFs compared to the next most competitive evaluation frameworks. More importantly, our Parallel runtime, illustrated in Figure 5, is significantly more efficient than competing frameworks. These empirical results validate our framework as a highly efficient tool for LALM evaluation.